Problem 6-56 (Static) Two-Stage Allocation and Product Costing (LO 6-5)

Owl-Eye Radiologists (OR) does various types of diagnostic imaging. Radiologists perform tests using sophisticated equipment. OR's management wants to compute the costs of performing tests for two different types of patients: those who are hospitalized (including those in emergency rooms) and those who are not hospitalized but are referred by physicians. The data for June for the two categories of patients follow.

Hospital Patients Other Patients Total
Units (i.e., procedures) produced 640 860 1,500
Equipment-hours used 240 120 360
Direct labor-hours 480 180 660
Direct labor costs $ 38,400 $ 10,800 $ 49,200
Overhead costs $ 49,560

The accountant first assigns overhead costs to two pools: overhead related to equipment-hours and overhead related to labor-hours. The analysis of overhead accounts by the cost accountant follows.

Account Amount Related to
Utilities $ 4,800 Equipment-hours
Supplies 12,600 Labor-hours
Indirect labor and supervision 20,400 Labor-hours
Equipment depreciation and maintenance 8,400 Equipment-hours
Miscellaneous 3,360 Equipment-hours


Required:

a. Compute the predetermined overhead rates assuming that Owl-Eye Radiologists uses equipment-hours to allocate equipment-related overhead costs and labor-hours to allocate labor-related overhead costs.

b. Compute the total costs of production and the cost per unit for each of the two types of patients undergoing tests in June.
Answer #1

a) Predetermine overhead rate :

Equipment related overhead cost = (4800+8400+3360)/360 = 46 per equipment hour

Labor hour related overhead cost = (12600+20400)/660 = 50 per labor hour

b) Calculate total production cost and cost per unit

Hospital patients Other patients
Direct labor 38400 10800
Equipment related overhead 11040 5520
Direct labor hour related overhead 24000 9000
Total cost of production 73440 25320
Unit 640 860
Unit cost 114.75 29.44